Circular No. 8248

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 UD_{16} (LONEOS) An apparently asteroidal object with not-unusual motion, found by

LONEOS on Oct. 16 (the discovery observation given below, together with other astrometry, appeared on MPS 88336, 90581, and 91035 with the designation 2003 UD_{16}; initial orbit on MPO 53844), has been found by C. W. Hergenrother to show a circular, condensed 11′′ coma and no tail on coadded 900-s R-band CCD exposures taken on Nov. 30 with the Mt. Hopkins 1.2-m reflector (astrometry below measured by T. B. Spahr). Additional unpublished LINEAR observations also appear below.

2003 UT α_{2000} δ_{2000} Mag. Observer
Oct. 16.39886 2^{h}59^{m}44.^{s}23 +7^{o}21′24′′.7 18.8 LONEOS
Nov. 30.20174 2 29 20.86 +8 58 14.8 19.3 LINEAR
30.20642 2 29 20.72 +8 58 15.3  Hergenrother
30.21402 2 29 20.46 +8 58 17.0 18.5  _{”}
30.24099 2 29 19.45 +8 58 23.0 19.3 LINEAR
30.25389 2 29 19.05 +8 58 25.6 19.6  _{”}

The following improved orbital elements by B. G. Marsden are from 31 observations, Oct. 16–Nov. 30.

T = 2004 Mar. 3.6269 TT  ω = 5.^{o}0941
e = 0.381095  Ω = 55.1589 }2000.0
q = 3.650531 AU  i = 23.4248
a = 5.898372 AU n^{o} = 0.0688028 P = 14.325 years

NOVAE IN M31 Further to IAUC 8222, K. Hornoch reports the discovery of an appar-

ent nova located at α = 0^{h}42^{m}53^{s}.64, δ = +41^{o}18′45′′.9 (equinox 2000.0), which is 105′′ east and 157′′ north of the center of M31. Available R-band magnitudes: Nov. 2.878 UT, [18.6; 9.073, 16.9; 29.883, 18.7. The new object is also absent from many more of his images dating to 2002.

Hornoch also reports additional R-band magnitudes of two earlier apparent novae in M31, from images taken by P. Kušnirak and himself: (1) Hornoch’s nova on IAUC 8222: Oct. 14.480 UT, 17.8; 17.976, 18.2; Nov. 9.073, 18.4; 29.883, 19.9. (2) the nova announced on IAUC 8226: Oct. 1.990, [19.0; 17.976, 17.9.

2003 December 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green